OpenStudy (anonymous):

Can someone help with this?: Find two numbers who difference is 60 and whose product is a minimum. ______smaller value ______larger value

OpenStudy (amistre64):

x-y= 60. xy= min. How minimum is minimum? 60-0=60; and 60*0=0.